OpenStudy (anonymous):

Bobby is bringing his collection of dimes and quarters to the bank to deposit into his savings account. If d represents the number of dimes in Bobby’s collection, q represents the number of quarters, and the collection is worth $63.75, which of the following equations models the value of Bobby’s collection

OpenStudy (anonymous):

A. 63.75 = d + q B. 63.75 = 25d + 10q C. 63.75 = 10d + 25q D. 63.75 = 0.10d + 0.25q